OpenStudy (anonymous):

A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. @hamidi

OpenStudy (hamidi):

give me a min or two

OpenStudy (anonymous):

a)-3x+y=-5 b)3x+y=-5 c)(y-8)=3(x-1) d)(y+8)=3(x+1)

OpenStudy (anonymous):

choose all that apply

OpenStudy (anonymous):

You should try posting this in the correct math section. But I will still try and help you

OpenStudy (anonymous):

ok sorry thanks for trying to help though

OpenStudy (anonymous):

Its fine

OpenStudy (hamidi):

i think its c

OpenStudy (anonymous):

any others?

OpenStudy (hamidi):

i dont believe

OpenStudy (anonymous):

ok can you help with more they will be under algebra

OpenStudy (hamidi):

ill try

OpenStudy (anonymous):

thanks

OpenStudy (hamidi):

ok see u in algebra

OpenStudy (anonymous):

nvm got to go getting off the plane

OpenStudy (hamidi):

what plane

OpenStudy (hamidi):

kik me some time IB_SIDDS